OpenStudy (anonymous):

Solve the equation by the square root property. (x+4)^2=4

OpenStudy (freckles):

take square root of both sides

OpenStudy (freckles):

keep in mind \[\sqrt{x^2}=|x|=\pm x\]

OpenStudy (anonymous):

Would it be -2,2?

OpenStudy (freckles):

that is if you x^2=9 then x=+/- sqrt(9) x=+/- 3

OpenStudy (freckles):

so you have x+4=?

OpenStudy (freckles):

right on

OpenStudy (freckles):

so solve x+4=2 and also solve x+4=-2

OpenStudy (anonymous):

so then -2,-6?

OpenStudy (freckles):

x=2-4=-2 check mark x=-2-4=-6 check mark yep

OpenStudy (anonymous):

thank you!